Dear all,
please find below a short summary of the activities related to the AO system.


People at LBTO:


R. Briguglio


remote support
M.Xompero, E.Pinna(VNC+Skype)


Summary
Positive environmental conditions allowed telescope activities tonight.
The shell was successfully set at 01:28UT. A preliminary inspection of the gap showed it is on its nominal range of values (24um<pos<39um).
The first part of the night was devoted to measure the AdSec astigmatism at changing elevation, also to improve the previoulsy calibrated LookUpTable. 14 points were sampled, with elevations ranging from 20° to 85°. At any position, a command was given to the shell and through IRTC the resulting image was observed. The correction value that resulted in the best image was taken as the astigmatism correction for that elevation. IRTC was operated remotely by J.Hill, who also took care of the inspection of the images. The result in shown in the following plot.

astig.png
To perform the measurement, the automatic high order offload was disabled: an error was found while re-enabling the automatic correction (forces are increasing instead of decreasing). This issue is under analysis now: please note that the correction worked fine during the previous days.

The second part of the night was devoted to WFS testing. The telescope was moved to NSV6945, an F5V, B=4.7 star, and images were collected both on the WFS technical viewer and he CCD39. Please find below the report of the W-Team that worked remotely.

The dome was closed at 08:51UT, because of wind speed and increasing cloud coverage.

W-Report follows:

A magnitude 4.7 star was pointed and the WFS unit was setup using the
parameters found during the alignment in the LBT clean room. Starlight
was immediately found on the pyramid and formed pupil images on the
ccd39 without the need of a star acquisition procedure. As expected,
atmospheric turbulence is prominent on single images. A sequence of 2000
frames (corresponding to 10 seconds of integration at the selected rate
of 200 Hz) at high modulation was acquired and averaged. The averaged
image clearly shows the central obstruction and swing arm shadow. The
features on the pupil borders are caused by the actuators removed from
the control loop in the last several days. Intensity is about 2000
counts/pixel, which is about as expected for this star magnitude and
acquisition rate.

Another test was attempted to refine the pupil position towards the
correct one for closed loop operation, but it was interrupted by the
onset of clouds.

Below: a screenshot of a single frame showing the turbolence (0.005
sec exposure), and the average of 2000 such frames.

sky_pupils.png


sky_pupils_average.png


Ciao and happy Easter to everybody

Runa

-- ArmandoRiccardi - 05 Apr 2010
I Attachment Action Size Date Who Comment
astig.pngpng astig.png manage 2 K 05 Apr 2010 - 13:55 ArmandoRiccardi Astigmatism vs elevation
sky_pupils.pngpng sky_pupils.png manage 10 K 05 Apr 2010 - 13:52 ArmandoRiccardi Pupils on sky (single shot)
sky_pupils_average.pngpng sky_pupils_average.png manage 9 K 05 Apr 2010 - 13:52 ArmandoRiccardi pupils on sky (averaging turbulnce)
Topic revision: r1 - 05 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback